OpenStudy (anonymous):

find the function this is finally graphed after the following transformations are applied to the graph y=sqrt x in the order listed 1. shift 7 units up 2. reflect about the x axis 3. reflect about the y axis

OpenStudy (anonymous):

\[f(x)\] \[f(x)+7\] \[-(f(x)+7)=-f(x)-7\] \[f(-x)-7\]

OpenStudy (anonymous):

ok but remember its not f(x) its y= 

OpenStudy (anonymous):

so the final graph would be sqrt x-7?

OpenStudy (anonymous):

take the steps i wrote, replace \[f(x)\] by \[\sqrt{x}\]

OpenStudy (anonymous):

Yes, you should know this applies to any f(x), not just sqrt(x). It really helps graphing if you understand transformations

OpenStudy (anonymous):

yea the answer is y= -sqrt -x -7